# Revealing the impact of synthetic native samples and multi-tasking strategies in Hindi-English code-mixed humour and sarcasm detection

### v1

@@ 4.1 Baseline results:
Llama-3.1 0.21 (0.37) $\underline{0.32}^{\#}$ 0.22
GPT-4 (0.78) 0.16 0.25 $\underline{0.42}^{\#}$
Table 4: Results of our experiment evaluating the impact of in-context learning on VMLMs. Notation: NLD for native language dataset, CM for code-mixed.
  • When we prompted the VMLMs with code-mixed few shots, Llama-3.1 achieved the highest F1-score (0.75) for humor detection, followed by GPT-4 (0.74) and Aya Expanse (0.73). In contrast, Gemma performed the worst with an F1-score of 0.09. For sarcasm detection, GPT-4 outperformed all models with an F1-score of 0.78, while others showed a sharp decline, i.e., Gemma (0.34), Aya Expanse (0.21), and Llama-3.1 (0.21).
  • On prompting with native humor few shots, Gemma model got confused and reported lower F1-scores ranging from 0.02 to 0.07. On contrary to this, Gemma model improved significantly when prompted on native sarcasm samples, the largest upto 0.47 F1-score, compared to 0.34 on code-mixed few shots case.
  • Similarly, Aya Expanse could not improve when prompted with native humor few shots. It maintained the F1-score in the range 0.67 - 0.73. But during prompting with native sarcasm few shots, it improved to 0.26 in terms of F1-score, compared to 0.21 when prompted with code-mixed few shots.
  • When prompted with native humor few shots, Llama-3.1 performed worse with the range of F1-score in 0.55 - 0.73. On the other hand, the native sarcasm few shots improved its F1-score upto 0.37, compared to 0.21 when prompted on code-mixed few shots.
  • Finally, GPT-4 also reported a sharp decrease in F1-score of both the tasks. In humor detection, the F1-scores ranged from 0.47 to 0.62, while in sarcasm detection, F1-scores ranged from 0.16 to 0.42.
  • We observed that none of the models could beat the best F1-scores observed while using MLMs. This insight aligned well with previous works Góes et al. (2023); Baranov et al. (2023); Zhang et al. (2024) which showed LLMs inability in detecting monolingual humor and sarcasm.
@@ 5 Conclusion:
  • Approach wise we got a F1-score of i) 0.79 for humor and 0.89 for sarcasm detection using native sample mixing, ii) 0.83 for humor and 0.91 for sarcasm detection using multitask learning, and iii) 0.75 for humor and 0.78 for sarcasm detection using in-context learning. Thus, among the three strategies, MTL performed best.
  • Among these three strategies, MTL reported the most significant improvement, with F1-score increments upto 0.08 for humor and 0.10 for sarcasm. Native sample mixing followed, with increments upto 0.05 for humor and 0.07 for sarcasm, while in-context learning showed no improvement in F1-scores.
@@ 6 Limitation:
  • Our MTL models occasionally failed on non-humorous and non-sarcastic samples. The possible reason behind it could be task and domain interference. We reported a detailed qualitative analysis of correctly and incorrectly classified samples in error analysis section (refer Appendix F.2).
  • Non-availability of native Hindi language datasets for these two tasks restricted us from utilizing raw Hindi samples for mixing in our experiments.
  • Due to the limited linguistic expertise, we evaluated our hypothesis only on the Hindi-English code-mixed scenarios.
  • We couldn’t make use of more larger VMLMs due to computational constraints. The larger ones may generalize the results more clearly.
As an extension of our work, we list down the future scopes as,
  • Other language pairs can be utilized to shed some light on the generalization of our approach to more languages.
  • Synthetic data generation for monolingual native languages can be another possible way to be evaluated, in the same lines.
  • Also, near native language samples can be tested on the similar lines, where we analyze the impact of mixing closely related languages.
@@
  • The pretraining of VMLMs should be more language inclusive, i.e., it should contain better data representation for code-mixed setting Zhang et al. (2023).
@@
Acknowledgment
The authors thank Rada Mihalcea and Diyi Yang for sharing their datasets.
@@ B.3.2 Native language datasets:
  • HASOC-20198 (HASOC): It contains 5,852 social media posts collected from Twitter and Facebook using hashtags and keywords. Following this, each sample was annotated as ‘hate’ or ‘non-hate’ by organizers of the HASOC track.
@@ C.2 Sarcasm:
Appendix D Experimental configurations:
@@ D.1 Experimental set-up:
Input: Text input tokens ($input$)
Given: BERT encoder ($BERT$), Taskwise last four layers of BERT module ($module_{task_{1}},module_{task_{2}}$), Gating scheme ($gate$)
@@
1:function MultiTaskModel($input$)
@@
4:     for $layer$ in $module_{task_{1}}$ do
@@
7:     for $layer$ in $module_{task_{2}}$ do
@@
The term $\|\mathbf{W}_{task_{1}}-\mathbf{W}_{task_{2}}\|_{2}$ represents the $L_{2}$-norm (Euclidean distance) between the weight matrices of the two tasks’ specific layers, enforcing soft parameter sharing between tasks.
@@
D.2 Model configurations:
We conducted all our experiments on a single NVIDIA A100 GPU card. We presented our considered set of VMLMs and their respective versions in Table 7. This set includes both open-source and closed-source source VMLMs. For closed source models like GPT-4, it’s important to note that their weights might change in the future as they are updated and improved. We conducted all our experiments with these models during the period from September, 2024 to November, 2024. We used a default set of hyperparameters for the VMLMs across runs to maintain consistency in results.
@@ D.2 Model configurations:
Appendix E Prompting details:
In this section, we described the prompting details for in-context learning for conducting Exp. 3 (refer Section 3.4). Our prompt consisted of three parts, i) system prompt, where we explained the task, ii) few shots, examples that we fed to learn from, and iii) user input, which is the query for which we needed the predicted label. We selected the few shots through clustering technique which is considered to be a better approach Huzaifah et al. (2024) than randomly picking examples. We presented the prompt template in Figure 4.
Figure 4: Prompt template for k-shot prompting utilized for humor detection.
Appendix F Ablation study & error analysis:

### v2

@@ 4.1 Baseline results:
Llama-3.1 0.21 $\textbf{0.45}^{\#}$ (0.51) 0.28
GPT-4 (0.78) 0.17 0.25 $0.43^{\#}$
Table 4: Results of our experiment evaluating the impact of in-context learning on VMLMs. Notation: CM for code-mixed.
                 Humor
  Model CM CM+Col CM+POTD CM+HaHa CM+16000
Gemma (0.77) 0.58 (0.77) 0.23 0.76
Aya Expanse 0.74 0.74 0.75 0.74 (0.76)
Llama-3.1 (0.77) 0.74 0.75 0.74 0.75
                 Sarcasm
  Model CM CM+NHD CM+iSarc CM+SC-V2
Gemma (0.74) 0.63 0.67 0.64
Aya Expanse 0.78 0.68 (0.79) 0.78
Llama-3.1 0.80 0.49 0.70 (0.81)
Table 5: Results of our experiment evaluating the impact of native language mixing in instruction fine-tuning of VMLMs using LoRA adapter. Notation: CM for code-mixed.
In this section, we reported our observations from the third experiment. The best F1-scores obtained from prompting VMLMs are reported in Table 4. A more detailed overview of obtained F1-scores is presented in Table LABEL:tab:results_vmlm_all_fewshot. The F1-scores obtained from instruction finetuning VMLMs are presented in Table 5. The highest overall scores for both tasks are marked in blue. Following were our takeaways,
  • When we prompted the VMLMs with code-mixed few shots, Llama-3.1 achieved the highest F1-score (0.75) for humour detection, followed by GPT-4 (0.74) and Aya Expanse (0.73). In contrast, Gemma performed the worst with an F1-score of 0.09. For sarcasm detection, GPT-4 outperformed all models with an F1-score of 0.78, while others showed a sharp decline, i.e., Gemma (0.34), Aya Expanse (0.21), and Llama-3.1 (0.21).
  • When prompted with native humour few-shot examples, Llama-3.1 maintained a stable F1-score between 0.74–0.75, while Aya Expanse showed no significant improvement, maintaining scores in the 0.71–0.74 range. In contrast, Gemma performed poorly, with F1-scores ranging from 0.07–0.11. GPT-4 also experienced a decline in performance, with F1-scores dropping to 0.56–0.62. Native few-shot prompting led to some improvements in sarcasm detection across models. Gemma’s performance significantly increased to 0.47, compared to 0.34 with code-mixed few-shots. Similarly, Aya Expanse improved to 0.26, up from 0.21, while Llama-3.1 achieved an F1-score of 0.51, a substantial increase from 0.21 in the code-mixed setting. However, GPT-4 continued to struggle, with F1-scores ranging from 0.17–0.43.
  • When VMLMs were finetuned using the code-mixed training set, Gemma and Llama-3.1 achieved the best F1-score of 0.77 in humour detection, while Llama-3.1 got best F1-score of 0.80 in sarcasm detection. When finetuned using native sample mixing strategy, Aya Expanse and Llama-3.1 maintained comparable performances within the range 0.74-0.76, whereas Gemma showed greater fluctuations (0.23-0.77). For sarcasm detection also, VMLMs showed varying fluctuations, i.e., Gemma (0.63-0.67), Aya Expanse (0.68-0.79), and Llama-3.1 (0.49-0.81). Although this finetuning approach outperformed the few-shot prompting method, it did not exceed the best F1-scores obtained with MLMs. Futhermore, a consistent pattern emerged from the native dataset point of view, where low KL divergence native datasets proved to be more effective for training compared to others.
  • The precision and recall scores showed a similar pattern. The exact values are reported in Appendix Table 13 and Table 14.
5 Error analysis:
  • For the ironic humor in Sl. No. 1, the humor arises from the switch between the political promise (’no power cuts’) and the ironic consequence (’high bills’) of the situation. Most of the models failed on it, except $\text{XLM-R}_{MTL}$ and $\text{MuRIL}_{MTL}$ as it had source of knowledge from other tasks like sarcasm.
@@ 6 Conclusion:
  • Among these three strategies, MTL reported the most significant improvement, with F1-score increments upto 10.67% for humor and 12.35% for sarcasm. Native sample mixing followed, with increments upto 6.76% for humor and 8.64% for sarcasm, while VMLMs in both set-up showed no improvement in F1-scores.
@@
  • The error analysis (refer Section 5) presented samples which justified the utility of related tasks in the multitask scenario.
  • The VMLMs showed poor performance due to a tendency to favor specific labels in sentiment-related classification tasks (refer Appendix D.3 for detailed observations). This insight aligned well with previous works Góes et al. (2023); Baranov et al. (2023); Zhang et al. (2024) which showed LLMs inability in detecting monolingual humor and sarcasm.
@@ 7 Limitation:
  • Our MTL models occasionally failed on non-humorous and non-sarcastic samples. The possible reason behind it could be task and domain interference. We reported a detailed qualitative analysis of correctly and incorrectly classified samples in error analysis section (refer Appendix 5).
  • Non-availability of native Hindi language datasets for these two tasks restricted us from utilizing gold labeled Hindi samples for mixing in our experiments.
  • Due to the limited linguistic expertise, we evaluated our hypothesis only on the Hindi-English code-mixed scenarios. Other language pairs can be utilized to shed some light on the generalization of our approach to more languages.
  • We couldn’t make use of more larger VMLMs due to computational constraints. The larger ones (with more than 100B parameters) may generalize the results more clearly.
  • Here, in our experiments we utilized the most widely used translator (Google Translate API) to generate synthetic Hindi samples. One can try other possible ways for synthetic data generation using various VMLMs, in the same lines as future scope.
  • We explored the impact of native samples for code-mixed humor and sarcasm detection. As future scope, one can test the impact of near native language (languages which have similar origin) samples and more closely associated tasks as well.
@@
  • The pretraining of VMLMs can be more language inclusive, i.e., it should contain better data representation for code-mixed setting Zhang et al. (2023).
@@ A.7 Research gap:
A.8 Code-mixing:
Code-mixing is a linguistic phenomenon common across multilingual speakers. Multilingual speakers are believed to outnumber monolingual speakers globally Tucker (1999). A large chunck of population in Asia, Europe, and North America know more then one language, this allows them to communicate among themselves by switching languages in a single utterance. It is more prominent in informal communications like social media posts and voice mails Patro et al. (2017).
@@ B.3.2 Native language datasets:
  • HASOC-20199 (HASOC): It contains 5,852 social media posts collected from Twitter and Facebook using hashtags and keywords. Following this, each sample was annotated as ‘hate’ or ‘non-hate’ by organizers of the HASOC track.
@@ C.2 Sarcasm:
Appendix D Additional details of Exp. 3:
D.1 Prompting details:
In this section, we described the prompting details for in-context learning for conducting Exp. 3 (refer Section 3.4). Our prompt consisted of three parts, i) system prompt, where we explained the task, ii) few shots, examples that we fed to learn from, and iii) user input, which is the query for which we needed the predicted label. We selected the few shots through clustering technique which is considered to be a better approach Huzaifah et al. (2024) than randomly picking examples. We presented the prompt template in Figure 2. The detailed results for the first scenario of few-shot prompting VMLM (Experiment-3) is presented in Table-LABEL:tab:results_vmlm_all_fewshot. The VMLMs were prompted with 0-shot and k-shot examples given in the context. The user input query remained in code-mixed language in each of the cases.
Figure 2: Prompt template for k-shot prompting utilized for humor detection. English parts are marked in red and the Hindi parts are marked in blue.
D.2 LoRA-adapter based finetuning:
For the second scenario, we performed instruction finetuning on the VMLMs using the default parameters of LoRA adapter based supervised finetuning given in LLaMA-Factory Zheng et al. (2024). We utilized the same prompt template for giving instruction to the models.
D.3 Detailed observations from Exp. 3:
In this section, we reported the additional observations of third experiment. While investigating the poor performance of the VMLMs, we identified certain patterns. The VMLMs were prompted to provide reasoning for their label predictions. Here are some observations,
  • VMLMs favored specific labels like humour and sarcasm, where it extracted comedic or ironic effect in plain non-humourous statements. For instance, in the non-humourous sample So jao sab, kal Monday hai.’ (Gloss: ‘Go to sleep, everyone. Tomorrow is Monday.’), Llama-3.1 predicted it as humourous and stated the reason : ‘The input is a Hindi phrase that translates to “So go everyone, it’s Monday." The humour lies in the fact that it’s a common expression that people use to bid farewell on Fridays, but it’s being used on Monday, which is the opposite day. It’s a play on expectations and a lighthearted way to poke fun at the monotony of the workweek.’.
  • We could find instances where VMLM assumed some prior context over the original non-sarcastic sample Kabhi triple talaq halala pe v kuch post kardo agar himmat hai to (Gloss: ‘Try posting something about triple talaq and halala if you have the courage.’), GPT-4 predicted it as sarcastic and stated the reason as ‘The statement is sarcastic as it challenges the recipient’s courage (“agar himmat hai to"), implying reluctance or hypocrisy. It uses a mocking tone to suggest selective bias in posting about issues. The indirect criticism and taunting nature make it a clear case of sarcasm.’
  • VMLMs showed a tendency to pick keywords, like in the non-sarcastic example Yeh bhee cricket kee deewani hain….’ (Gloss: ‘She is also crazy about cricket…’), Llama-3.1 predicted it sarcastic and gave the reason as ‘The phrase “Yeh bhee cricket kee deewani hain" translates to "You are also a cricket fanatic" in English. However, the use of the word “bhee" (also) in a sentence suggests that the speaker is being sarcastic, implying that the person is extremely obsessed with cricket, which is often considered a stereotypical trait in Indian culture. The tone of the sentence suggests that the speaker is mocking or ridiculing the person’s enthusiasm for cricket.’
Appendix E Experimental configurations:
@@ E.1 Experimental set-up:
Input: Text input tokens (inputinput)
Given: BERT encoder ($BERT$), Taskwise last four layers of BERT module ($module_{task_{1}},module_{task_{2}}$), Gating scheme (gategate)
@@
1:function MultiTaskModel(inputinput)
@@
4:  for layerlayer in $module_{task_{1}}$ do
@@
7:  for layerlayer in $module_{task_{2}}$ do
@@
For each task $t\in\{1,2\}$, the final task-specific representation is computed by gating the shared BERT representation ($\mathbf{h}_{\text{BERT}}$) with the task-specific features ($\mathbf{h}_{\text{task}_{t}}$):
$\mathbf{o}_{t}$ $=\text{Gate}(\mathbf{h}_{\text{BERT}},\mathbf{h}_{\text{task}_{t}})$
$=\boldsymbol{\alpha}_{t}\odot\mathbf{h}_{\text{BERT}}+(1-\boldsymbol{\alpha}_{t})\odot\mathbf{h}_{\text{task}_{t}}$ (1)
where the gate coefficient $\boldsymbol{\alpha}_{t}$ is computed as:
$\boldsymbol{\alpha}_{t}$ $=\sigma\big{(}\mathbf{W}_{g,t}[\mathbf{h}_{\text{BERT}}\|\mathbf{h}_{\text{task}_{t}}]+\mathbf{b}_{g,t}\big{)}$ (2)
Here:
  • $\mathbf{h}_{\text{BERT}}$: Final hidden states from shared replica of BERT.
  • $\mathbf{h}_{\text{task}_{t}}$: Task-specific hidden representation after processing through $\text{module}_{task_{t}}$.
  • $\sigma$: Sigmoid activation $\in(0,1)$.
  • $\mathbf{W}_{g,t}\in\mathbb{R}^{D\times 2D},\mathbf{b}_{g,t}\in\mathbb{R}^{D}$: Gating parameters for task $t$.
@@
The term $\|\mathbf{W}_{task_{1}}-\mathbf{W}_{task_{2}}\|_{2}$ represents the $L_{2}$-norm (Euclidean distance) between the weight matrices of the two tasks’ specific layers, enforcing soft parameter sharing between tasks Rotman and Reichart (2019).
@@
E.2 Model configurations:
We conducted all our experiments on a single NVIDIA A100 GPU card. We presented our considered set of VMLMs and their respective versions in Table 8. This set includes both open-source and closed-source source VMLMs. For closed source models like GPT-4, it’s important to note that their weights might change in the future as they are updated and improved. We conducted all our experiments with these models during the period from September, 2024 to November, 2024. We used a default set of hyperparameters for VMLMs using LLaMA-Factory Zheng et al. (2024) across runs to maintain consistency in the results (see Table 9).
@@ E.2 Model configurations:
Parameter Value
do_sample true
temperature 0.95
top_p 0.7
top_k 50
num_beams 1
max_length 1024
max_new_tokens 1024
repetition_penalty 1.0
length_penalty 1.0
skip_special_tokens true
Table 9: Generation parameters in VMLMs.
Appendix F Ablation study:
@@ Appendix G Examination of translated Hindi data:
                 Humor
Precision
  NLD $\rightarrow$ Col POTD HaHa 16000
NB 0.72 0.78 0.74 0.77 0.79 0.74 0.79 0.78 0.75 0.79 0.76 0.78 0.73
RF 0.72 0.72 0.73 0.74 0.74 0.74 0.75 0.73 0.73 0.75 0.74 0.73 0.72
SVM 0.73 0.73 0.70 0.71 0.73 0.70 0.74 0.72 0.74 0.73 0.72 0.72 0.74
mBERT 0.78 0.77 0.75 0.71 0.74 0.80 0.77 0.73 0.76 0.78 0.78 0.79 0.73
XLM-R 0.73 0.75 0.75 0.74 0.77 0.77 0.75 0.76 0.78 0.77 0.76 0.78 0.76
MuRIL 0.77 0.76 0.76 0.77 0.75 0.75 0.80 0.78 0.81 0.78 0.77 0.78 0.80
IndicBERT 0.70 0.78 0.77 0.74 0.79 0.76 0.79 0.74 0.80 0.77 0.74 0.77 0.76
                 Recall
  NLD $\rightarrow$ Col POTD HaHa 16000
NB 0.78 0.72 0.76 0.73 0.71 0.76 0.71 0.70 0.74 0.71 0.74 0.72 0.79
RF 0.72 0.67 0.68 0.65 0.66 0.70 0.67 0.70 0.68 0.69 0.65 0.65 0.64
SVM 0.69 0.64 0.60 0.63 0.65 0.70 0.65 0.67 0.65 0.65 0.64 0.67 0.65
mBERT 0.78 0.70 0.70 0.61 0.69 0.72 0.73 0.64 0.76 0.74 0.70 0.71 0.68
XLM-R 0.77 0.70 0.71 0.72 0.69 0.73 0.71 0.72 0.73 0.70 0.76 0.74 0.68
MuRIL 0.73 0.83 0.68 0.75 0.71 0.83 0.76 0.78 0.73 0.78 0.76 0.77 0.76
IndicBERT 0.79 0.75 0.68 0.69 0.75 0.77 0.71 0.84 0.72 0.73 0.75 0.69 0.76
                 Sarcasm
Precision
  NLD $\rightarrow$ NHD iSarc SC-V2
NB 0.71 0.49 0.38 0.42 0.56 0.47 0.50 0.34 0.35 0.46
RF 0.84 0.50 0.60 0.57 0.52 0.56 0.58 0.53 0.61 0.79
SVM 0.73 0.57 0.60 0.76 0.62 0.63 0.71 0.82 0.76 0.72
mBERT 0.93 0.83 0.83 0.76 0.82 0.84 0.80 0.82 0.85 0.82
XLM-R 0.83 0.78 0.83 0.82 0.85 0.82 0.83 0.79 0.82 0.85
MuRIL 0.84 0.87 0.88 0.83 0.87 0.89 0.86 0.83 0.85 0.84
IndicBERT 0.82 0.86 0.85 0.86 0.81 0.88 0.89 0.82 0.84 0.83
                 Recall
  NLD $\rightarrow$ NHD iSarc SC-V2
NB 0.80 0.29 0.36 0.40 0.31 0.39 0.37 0.33 0.34 0.33
RF 0.59 0.38 0.47 0.63 0.51 0.59 0.55 0.44 0.60 0.50
SVM 0.79 0.61 0.59 0.72 0.67 0.66 0.71 0.64 0.63 0.73
mBERT 0.70 0.83 0.72 0.85 0.77 0.78 0.75 0.75 0.79 0.88
XLM-R 0.80 0.85 0.76 0.83 0.78 0.80 0.80 0.84 0.83 0.79
MuRIL 0.82 0.85 0.89 0.81 0.83 0.82 0.84 0.91 0.85 0.90
IndicBERT 0.80 0.86 0.87 0.84 0.81 0.88 0.78 0.84 0.82 0.85
Table 10: Precision and recall scores for humor (first and second table) and sarcasm (third and fourth table) detection using native sample mixing. Notation: NLD for native language dataset, CM for code-mixed. The scores of best-performing models for the individual training scenarios are marked in bold and highest in the respective task is marked in blue.
       Humor
  Precision Recall
NLD : Col $\text{mBERT}_{MTL}$ $\text{XLM-R}_{MTL}$ $\text{MuRIL}_{MTL}$ $\text{mBERT}_{MTL}$ $\text{XLM-R}_{MTL}$ $\text{MuRIL}_{MTL}$
Hate Sarcasm Gate w/o Gate Gate w/o Gate Gate w/o Gate Gate w/o Gate Gate w/o Gate Gate w/o Gate
[…]
0.74 0.76 0.76 0.77 0.77 0.77 0.79 0.72 0.71 0.75 0.77 0.75
[…]
NLD : POTD
[…]
0.77 0.76 0.79 0.77 0.75 0.73 0.68 0.80 0.74 0.80 0.67 0.70
[…]
NLD : HaHa
[…]
0.77 0.74 0.80 0.79 0.78 0.77 0.70 0.70 0.77 0.75 0.82 0.75
[…]
NLD : 16000
[…]
0.79 0.77 0.79 0.76 0.78 0.75 0.78 0.79 0.77 0.78 0.77 0.71
[…]
Table 11: Precision and recall scores of MTL framework and ablation study for humor detection. Notation: ‘NLD’ for Native Language Dataset. The scores of best-performing models for the individual training scenarios are marked in bold and highest in the respective task is marked in blue.
       Sarcasm
  Precision Recall
NLD : NHD $\text{mBERT}_{MTL}$ $\text{XLM-R}_{MTL}$ $\text{MuRIL}_{MTL}$ $\text{mBERT}_{MTL}$ $\text{XLM-R}_{MTL}$ $\text{MuRIL}_{MTL}$
Hate Humor Gate w/o Gate Gate w/o Gate Gate w/o Gate Gate w/o Gate Gate w/o Gate Gate w/o Gate
[…]
0.81 0.85 0.83 0.83 0.84 0.79 0.81 0.85 0.88 0.81 0.87 0.78
[…]
NLD : iSarc
[…]
0.82 0.81 0.82 0.83 0.85 0.76 0.81 0.85 0.83 0.79 0.85 0.73
[…]
NLD : SC-V2
[…]
0.83 0.83 0.84 0.83 0.87 0.81 0.82 0.83 0.87 0.80 0.82 0.81
[…]
Table 12: Precision and recall scores of MTL framework and ablation study for sarcasm detection. Notation: ‘NLD’ for Native Language Dataset. The scores of best-performing models for the individual training scenarios are marked in bold and highest in the respective task is marked in blue.
                 Humor
  Precision Recall
Model CM Col POTD HaHa 16000 CM Col POTD HaHa 16000
Gemma 0.06 0.05 0.05 0.06 0.06 0.66 0.66 0.66 0.67 0.66
Aya Expanse 0.63 0.64 0.63 0.60 0.63 0.85 0.85 0.85 0.84 0.85
Llama-3.1 0.61 0.62 0.65 0.59 0.62 0.94 0.86 0.86 0.86 0.86
GPT-4 0.65 0.46 0.51 0.45 0.44 0.86 0.80 0.81 0.79 0.79
                 Sarcasm
  Precision Recall
Model CM NHD iSarc SC-V2 CM NHD iSarc SC-V2
Gemma 0.26 0.08 0.29 0.37 0.74 0.67 0.76 0.78
Aya Expanse 0.12 0.12 0.15 0.13 0.94 0.94 0.93 0.93
Llama-3.1 0.12 0.31 0.31 0.18 0.94 0.78 0.80 0.70
GPT-4 0.65 0.12 0.14 0.39 0.94 0.67 0.94 0.76
Table 13: Precision and recall scores of our experiment evaluating the impact of in-context learning on VMLMs. Notation: CM for code-mixed. The scores of best-performing models for the individual training scenarios are marked in bold and highest in the respective task is marked in blue.
                 Humor
  Precision Recall
Model CM CM+Col CM+POTD CM+HaHa CM+16000 CM CM+Col CM+POTD CM+HaHa CM+16000
Gemma 0.71 0.52 0.64 0.35 0.62 0.87 0.78 0.96 0.66 0.94
Aya Expanse 0.60 0.66 0.67 0.66 0.62 0.93 0.84 0.87 0.86 0.94
Llama-3.1 0.65 0.66 0.67 0.66 0.67 0.96 0.84 0.85 0.84 0.85
                 Sarcasm
  Precision Recall
Model CM CM+NHD CM+iSarc CM+SC-V2 CM CM+NHD CM+iSarc CM+SC-V2
Gemma 0.60 0.59 0.60 0.58 0.93 0.80 0.81 0.80
Aya Expanse 0.66 0.52 0.66 0.66 0.93 0.94 0.95 0.93
Llama-3.1 0.67 0.42 0.54 0.69 0.96 0.75 0.93 0.96
Table 14: Precision and recall scores of our experiment evaluating the impact of native language mixing in instruction fine-tuning of VMLMs using LoRA adapter. Notation: CM for code-mixed. The scores of best-performing models for the individual training scenarios are marked in bold and highest in the respective task is marked in blue.